# Beam tracing and profile evolution for localised beams in inhomogeneous plasmas

### v2

@@ A.3.1 Terms proportional to $\bm{w}\bm{w}$
$\begin{split}&-\frac{1}{2}w_{k}w_{\rho}\Bigg(\hat{e}^{*}_{i}\partial K_{\nu}(D_{im})\hat{e}_{m}\frac{\hat{g}_{\nu}}{g}\left(\frac{\mathrm{d}\Psi_{w,k\rho}}{\mathrm{d}\tau}+2\frac{\mathrm{d}\hat{g}_{k}}{\mathrm{d}\tau}\frac{\mathrm{d}K_{\rho}}{\mathrm{d}\tau}\right)\\ &\qquad+\Psi_{k\mu}\Psi_{\rho\nu}\hat{e}^{*}_{i}\partial K_{\mu}\partial K_{\nu}(D_{im})\hat{e}_{m}+2\hat{e}^{*}_{i}\partial_{k}(D_{im})(\Psi_{\rho l}\partial K_{l}+\partial_{\rho})\hat{e}_{m}\\ &\qquad+\hat{e}^{*}_{i}\partial_{\mu}\partial_{\nu}(D_{im})\hat{e}_{m}+2\hat{e}^{*}_{i}\Psi_{k\nu}\partial K_{\nu}(D_{im})(\Psi_{\rho l}\partial K_{l}+\partial_{\rho})\hat{e}_{m}\Bigg)=0.\end{split}$ (136)
@@
Note that anything along $\bm{g}$ is contracted to 0 by $w_{k}w_{\rho}$. We can utilise this throughout the expression, but more usefully, we can notice that $\left(\frac{\mathrm{d}\Psi_{w,k\rho}}{\mathrm{d}\tau}+2\frac{\mathrm{d}\hat{g}_{(k}}{\mathrm{d}\tau}\frac{\mathrm{d}K_{\rho)}}{\mathrm{d}\tau}\right)$ can simply be replaced with $\frac{\mathrm{d}\Psi_{k\rho}}{\mathrm{d}\tau}$. This yields the following equation:
$\begin{split}&\frac{1}{2}w_{k}w_{\rho}\Bigg(\frac{\mathrm{d}\Psi_{k\rho}}{\mathrm{d}\tau}+\Psi_{k\mu}\Psi_{\rho\nu}\hat{e}^{*}_{i}\partial K_{\mu}\partial K_{\nu}(D_{im}\hat{e}_{m})+\hat{e}^{*}_{i}\partial_{k}\partial_{\rho}(D_{im}\hat{e}_{m})\\ &\qquad+\hat{e}^{*}_{i}\Psi_{k\nu}\partial K_{\nu}\partial_{\rho}(D_{im}\hat{e}_{m})+\hat{e}^{*}_{i}\Psi_{\rho\nu}\partial K_{\nu}\partial_{k}(D_{im}\hat{e}_{m})\Bigg)=0.\end{split}$ (138)
We then apply the second equality of eq.137, and then call upon the fact that $\Psi_{\mu\nu}\partial K_{\nu}H=-\partial_{\mu}H$ to simplify the above expression, obtaining:
$\begin{split}\frac{1}{2}w_{k}w_{\rho}\left(\frac{\mathrm{d}\Psi_{k\rho}}{\mathrm{d}\tau}+\Psi_{k\mu}\Psi_{\rho\nu}\partial K_{\mu}\partial K_{\nu}(H)+\partial_{k}\partial_{\rho}(H)+\Psi_{k\nu}\partial K_{\nu}\partial_{\rho}(H)+\Psi_{\rho\nu}\partial K_{\nu}\partial_{k}(H)\right)=0\end{split}$ (139)
We can also differentiate $\Psi_{\mu\nu}\partial K_{\nu}H=-\partial_{\mu}H$ with respect to $\tau$ to yield the components of the beam tracing equation with a $\bm{g}$ component.
$\begin{split}\frac{\mathrm{d}}{\mathrm{d}\tau}(\Psi_{\mu\nu}\partial K_{\nu}H+\partial_{\mu}H)&=0\\ g_{k}\left(\frac{\mathrm{d}\Psi_{k\rho}}{\mathrm{d}\tau}+\Psi_{k\mu}\Psi_{\rho\nu}\partial K_{\mu}\partial K_{\nu}(H)+\partial_{k}\partial_{\rho}(H)+\Psi_{k\nu}\partial K_{\nu}\partial_{\rho}(H)+\Psi_{\rho\nu}\partial K_{\nu}\partial_{k}(H)\right)&=0.\end{split}$ (140)
@@
$\begin{split}\frac{\mathrm{d}\Psi_{k\rho}}{\mathrm{d}\tau}+\Psi_{k\mu}\Psi_{\rho\nu}\partial K_{\mu}\partial K_{\nu}(H)+\partial_{k}\partial_{\rho}(H)+\Psi_{k\nu}\partial K_{\nu}\partial_{\rho}(H)+\Psi_{\rho\nu}\partial K_{\nu}\partial_{k}(H)=0.\end{split}$ (141)
These equations provide us with the evolution of the Gaussian envelope as it propagates through the plasma. More specifically, we have chosen our function, $\Psi_{\mu\nu}(\tau)$ to satisfy eq.18.
@@ A.3.2 Terms independent of $\bm{w}$
$\begin{split}\alpha^{(0)}\Bigg(&\frac{\mathrm{dln}(\alpha^{(0)})}{\mathrm{d}\tau}+\frac{\hat{g}_{\mu}}{g}\partial K_{\mu}D_{im}\hat{e}^{*}_{i}\frac{\mathrm{d}\hat{e}_{m}}{\mathrm{d}\tau}+\hat{e}^{*}_{i}\partial K_{\mu}D_{im}(\Psi_{\mu\nu}\partial K_{\nu}+\partial_{\mu})\hat{e}_{m}\\ &-\hat{g}_{k}\partial K_{k}D_{im}\hat{e}^{*}_{i}\hat{g}_{\mu}(\Psi_{\mu\nu}\partial K_{\nu}+\partial_{\mu})\hat{e}_{m}+\frac{1}{2}\Psi_{\mu\nu}\hat{e}^{*}_{i}\partial K_{\mu}\partial K_{\nu}D_{im}\hat{e}_{m}\Bigg)=0.\end{split}$ (142)
Using the fact that $g_{\mu}\Psi_{\mu\nu}=\frac{\mathrm{d}K_{\nu}}{\mathrm{d}\tau}$, and that $g_{\mu}\partial_{\mu}+\frac{\mathrm{d}K_{\mu}}{\mathrm{d}\tau}\partial K_{\mu}=\frac{\mathrm{d}}{\mathrm{d}\tau}$, we can then simplify into the following equation:
$\begin{split}\frac{\mathrm{dln}(\alpha^{(0)})}{\mathrm{d}\tau}+\frac{\mathrm{d}K_{\mu}}{\mathrm{d}\tau}\hat{e}^{*}_{m}\partial K_{\mu}\hat{e}_{m}+\hat{e}^{*}_{i}\partial K_{\mu}D_{im}\partial_{\mu}\hat{e}_{m}+\frac{1}{2}\Psi_{\mu\nu}\partial K_{\mu}\partial K_{\nu}H=0.\end{split}$ (143)
With this, we then decompose the amplitude into its real and imaginary parts, since $\ln(\alpha^{(0)})=\ln|\alpha^{(0)}|+\mathrm{i}\phi$ and analyse each separately. Using the fact that:
$\partial_{\mu}\partial K_{\mu}(\hat{e}^{*}_{i}D_{im}\hat{e}_{m})=\hat{e}^{*}_{i}\partial K_{\mu}D_{im}\partial_{\mu}\hat{e}_{m}+\partial_{\mu}\hat{e}^{*}_{i}\partial K_{\mu}D_{im}\hat{e}_{m}.$ (144)
@@
$\frac{\mathrm{dln}(|\alpha^{(0)}|)}{\mathrm{d}\tau}+\frac{1}{2}\partial K_{\mu}\partial_{\mu}H+\frac{1}{2}\Re(\Psi_{\mu\nu})\partial K_{\mu}\partial K_{\nu}H=0.$ (145)
@@
$\frac{\mathrm{dln}(|\alpha^{(0)}|)}{\mathrm{d}\tau}-\frac{1}{4}\left(\frac{\mathrm{dln}\left(\mathrm{Det(}\Im(\Psi)_{\mu\nu})\right)}{\mathrm{d}\tau}-2\hat{g}_{\mu}\hat{g}_{\nu}(\partial K_{\mu}\partial_{\nu}H+\Re(\Psi_{\mu k})\partial K_{k}\partial K_{\nu}H)\right)=0.$ (149)
@@
$\frac{\mathrm{dln}(|\alpha^{(0)}|)}{\mathrm{d}\tau}=\frac{1}{4}\frac{\mathrm{dln}\left(\mathrm{Det(}\Im(\Psi)_{\mu\nu})\right)}{\mathrm{d}\tau}-\frac{1}{2g}\frac{\mathrm{d}g}{\mathrm{d}\tau},$ (150)
@@
$|\alpha^{(0)}|=C\frac{\mathrm{Det(}\Im(\Psi)_{\mu\nu})^{\frac{1}{4}}}{g^{\frac{1}{2}}}.$ (151)
@@
$\begin{split}&\frac{\mathrm{d}\phi}{\mathrm{d}\tau}+\frac{1}{\mathrm{i}}\frac{\mathrm{d}K_{\mu}}{\mathrm{d}\tau}\hat{e}^{*}_{m}\partial K_{\mu}\hat{e}_{m}+\frac{1}{2}\Im(\Psi_{\mu\nu})\partial K_{\mu}\partial K_{\nu}H\\ &+\frac{1}{2\mathrm{i}}(\hat{e}^{*}_{i}\partial K_{\mu}D_{im}\partial_{\mu}\hat{e}_{m}-\partial_{\mu}\hat{e}^{*}_{i}\partial K_{\mu}D_{im}\hat{e}_{m})=0.\end{split}$ (152)
From here we just use the fact that $\partial K_{\mu}H\hat{e}_{i}=\partial K_{\mu}D_{im}\hat{e}_{m}+D_{im}\partial K_{\mu}\hat{e}_{m}$ along with its conjugate equation to arrive at this final simple expression for the phase evolution:
$\frac{\mathrm{d}\phi}{\mathrm{d}\tau}=\mathrm{i}\hat{e}^{*}_{i}\frac{\mathrm{d}\hat{e}_{i}}{\mathrm{d}\tau}+\frac{\mathrm{i}}{2}(\partial_{\mu}\hat{e}^{*}_{i}D_{im}\partial K_{\mu}\hat{e}_{m}-\partial K_{\mu}\hat{e}^{*}_{i}D_{im}\partial_{\mu}\hat{e}_{m})-\frac{1}{2}\Im(\Psi_{\mu\nu})\partial K_{\mu}\partial K_{\nu}H.$ (153)
@@
$\frac{\mathrm{d}\phi_{P}}{\mathrm{d}\tau}=\mathrm{i}\hat{e}^{*}_{i}\frac{\mathrm{d}\hat{e}_{i}}{\mathrm{d}\tau}+\frac{\mathrm{i}}{2}(\partial_{\mu}\hat{e}^{*}_{i}D_{im}\partial K_{\mu}\hat{e}_{m}-\partial K_{\mu}\hat{e}^{*}_{i}D_{im}\partial_{\mu}\hat{e}_{m}).$ (154)
@@
$\frac{\mathrm{d}\phi_{G}}{\mathrm{d}\tau}=-\frac{1}{2}\Im(\Psi_{\mu\nu})\partial K_{\mu}\partial K_{\nu}H.$ (155)
We have thus successfully defined $\alpha^{(0)}(\tau)$ and $\Psi_{\mu\nu}(\tau)$ such that our terms that do not involve derivatives of our profile function are set to 0.

### v3

@@ A.3.1 Terms proportional to $\bm{w}\bm{w}$
$\begin{split}&-\frac{1}{2}w_{k}w_{\rho}\Bigg(\hat{e}^{*}_{i}\frac{\partial}{\partial K_{\nu}}(D_{im})\hat{e}_{m}\frac{\hat{g}_{\nu}}{g}\left(\frac{\mathrm{d}\Psi_{w,k\rho}}{\mathrm{d}\tau}+2\frac{\mathrm{d}\hat{g}_{k}}{\mathrm{d}\tau}\frac{\mathrm{d}K_{\rho}}{\mathrm{d}\tau}\right)\\ &\qquad+\Psi_{k\mu}\Psi_{\rho\nu}\hat{e}^{*}_{i}\frac{\partial}{\partial K_{\mu}}\frac{\partial}{\partial K_{\nu}}(D_{im})\hat{e}_{m}+2\hat{e}^{*}_{i}\partial_{k}(D_{im})(\Psi_{\rho l}\frac{\partial}{\partial K_{l}}+\partial_{\rho})\hat{e}_{m}\\ &\qquad+\hat{e}^{*}_{i}\partial_{\mu}\partial_{\nu}(D_{im})\hat{e}_{m}+2\hat{e}^{*}_{i}\Psi_{k\nu}\frac{\partial}{\partial K_{\nu}}(D_{im})(\Psi_{\rho l}\frac{\partial}{\partial K_{l}}+\partial_{\rho})\hat{e}_{m}\Bigg)=0.\end{split}$ (127)
@@
Note that anything along $\bm{g}$ is contracted to 0 by $w_{k}w_{\rho}$. We can utilise this throughout the expression, but more usefully, we can notice that:
$\left(\frac{\mathrm{d}\Psi_{w,\mu\nu}}{\mathrm{d}\tau}+\frac{\mathrm{d}\hat{g}_{\mu}}{\mathrm{d}\tau}\frac{\mathrm{d}K_{\nu}}{\mathrm{d}\tau}+\frac{\mathrm{d}\hat{g}_{\nu}}{\mathrm{d}\tau}\frac{\mathrm{d}K_{\mu}}{\mathrm{d}\tau}\right)_{w}=\left(\frac{\mathrm{d}\Psi_{k\rho}}{\mathrm{d}\tau}\right)_{w}.$ (129)
This yields the following equation:
$\begin{split}&\frac{1}{2}w_{k}w_{\rho}\Bigg(\frac{\mathrm{d}\Psi_{k\rho}}{\mathrm{d}\tau}+\Psi_{k\mu}\Psi_{\rho\nu}\hat{e}^{*}_{i}\frac{\partial}{\partial K_{\mu}}\frac{\partial}{\partial K_{\nu}}(D_{im}\hat{e}_{m})+\hat{e}^{*}_{i}\partial_{k}\partial_{\rho}(D_{im}\hat{e}_{m})\\ &\qquad+\hat{e}^{*}_{i}\Psi_{k\nu}\frac{\partial}{\partial K_{\nu}}\partial_{\rho}(D_{im}\hat{e}_{m})+\hat{e}^{*}_{i}\Psi_{\rho\nu}\frac{\partial}{\partial K_{\nu}}\partial_{k}(D_{im}\hat{e}_{m})\Bigg)=0.\end{split}$ (130)
We then apply the second equality of eq.128, and then call upon the fact that $\Psi_{\mu\nu}\frac{\partial}{\partial K_{\nu}}H=-\partial_{\mu}H$ to simplify the above expression, obtaining:
$\begin{split}\frac{1}{2}w_{k}w_{\rho}\left(\frac{\mathrm{d}\Psi_{k\rho}}{\mathrm{d}\tau}+\Psi_{k\mu}\Psi_{\rho\nu}\frac{\partial}{\partial K_{\mu}}\frac{\partial}{\partial K_{\nu}}(H)+\partial_{k}\partial_{\rho}(H)+\Psi_{k\nu}\frac{\partial}{\partial K_{\nu}}\partial_{\rho}(H)+\Psi_{\rho\nu}\frac{\partial}{\partial K_{\nu}}\partial_{k}(H)\right)=0.\end{split}$ (131)
We can also differentiate $\Psi_{\mu\nu}\frac{\partial}{\partial K_{\nu}}H=-\partial_{\mu}H$ with respect to $\tau$ to yield the components of the beam tracing equation with a $\bm{g}$ component.
$\begin{split}\frac{\mathrm{d}}{\mathrm{d}\tau}(\Psi_{\mu\nu}\frac{\partial}{\partial K_{\nu}}H+\partial_{\mu}H)&=0,\\ g_{k}\left(\frac{\mathrm{d}\Psi_{k\rho}}{\mathrm{d}\tau}+\Psi_{k\mu}\Psi_{\rho\nu}\frac{\partial}{\partial K_{\mu}}\frac{\partial}{\partial K_{\nu}}(H)+\partial_{k}\partial_{\rho}(H)+\Psi_{k\nu}\frac{\partial}{\partial K_{\nu}}\partial_{\rho}(H)+\Psi_{\rho\nu}\frac{\partial}{\partial K_{\nu}}\partial_{k}(H)\right)&=0.\end{split}$ (132)
@@
$\begin{split}\frac{\mathrm{d}\Psi_{k\rho}}{\mathrm{d}\tau}+\Psi_{k\mu}\Psi_{\rho\nu}\frac{\partial}{\partial K_{\mu}}\frac{\partial}{\partial K_{\nu}}(H)+\partial_{k}\partial_{\rho}(H)+\Psi_{k\nu}\frac{\partial}{\partial K_{\nu}}\partial_{\rho}(H)+\Psi_{\rho\nu}\frac{\partial}{\partial K_{\nu}}\partial_{k}(H)=0.\end{split}$ (133)
These equations provide us with the evolution of the Gaussian envelope as it propagates through the plasma. More precisely, we have chosen our tensor, $\bm{\Psi}$, to satisfy eq.19, as such a tensor trivially satisfies eq. 133.
@@ A.3.2 Terms independent of $\bm{w}$
$\begin{split}A^{(0)}\Bigg(&\frac{\mathrm{dln}(A^{(0)})}{\mathrm{d}\tau}+\frac{\hat{g}_{\mu}}{g}\frac{\partial}{\partial K_{\mu}}D_{im}\hat{e}^{*}_{i}\frac{\mathrm{d}\hat{e}_{m}}{\mathrm{d}\tau}+\hat{e}^{*}_{i}\frac{\partial}{\partial K_{\mu}}D_{im}\left(\Psi_{\mu\nu}\frac{\partial}{\partial K_{\nu}}+\partial_{\mu}\right)\hat{e}_{m}\\ &-\hat{g}_{k}\partial K_{k}D_{im}\hat{e}^{*}_{i}\hat{g}_{\mu}\left(\Psi_{\mu\nu}\frac{\partial}{\partial K_{\nu}}+\partial_{\mu}\right)\hat{e}_{m}+\frac{1}{2}\Psi_{\mu\nu}\hat{e}^{*}_{i}\frac{\partial}{\partial K_{\mu}}\frac{\partial}{\partial K_{\nu}}D_{im}\hat{e}_{m}\Bigg)=0.\end{split}$ (134)
Using the fact that $g_{\mu}\Psi_{\mu\nu}=\frac{\mathrm{d}K_{\nu}}{\mathrm{d}\tau}$, and that $g_{\mu}\partial_{\mu}+\frac{\mathrm{d}K_{\mu}}{\mathrm{d}\tau}\frac{\partial}{\partial K_{\mu}}=\frac{\mathrm{d}}{\mathrm{d}\tau}$, we can then simplify into the following equation:
$\begin{split}\frac{\mathrm{dln}(A^{(0)})}{\mathrm{d}\tau}+\frac{\mathrm{d}K_{\mu}}{\mathrm{d}\tau}\hat{e}^{*}_{m}\frac{\partial}{\partial K_{\mu}}\hat{e}_{m}+\hat{e}^{*}_{i}\frac{\partial}{\partial K_{\mu}}D_{im}\partial_{\mu}\hat{e}_{m}+\frac{1}{2}\Psi_{\mu\nu}\frac{\partial}{\partial K_{\mu}}\frac{\partial}{\partial K_{\nu}}H=0.\end{split}$ (135)
With this, we then decompose the amplitude into its real and imaginary parts, since $\ln(A^{(0)})=\ln|A^{(0)}|+\mathrm{i}\phi$ and analyse each separately. Using the fact that:
$\partial_{\mu}\frac{\partial}{\partial K_{\mu}}(\hat{e}^{*}_{i}D_{im}\hat{e}_{m})=\hat{e}^{*}_{i}\frac{\partial}{\partial K_{\mu}}D_{im}\partial_{\mu}\hat{e}_{m}+\partial_{\mu}\hat{e}^{*}_{i}\frac{\partial}{\partial K_{\mu}}D_{im}\hat{e}_{m}.$ (136)
@@
$\frac{\mathrm{dln}(|A^{(0)}|)}{\mathrm{d}\tau}+\frac{1}{2}\frac{\partial}{\partial K_{\mu}}\partial_{\mu}H+\frac{1}{2}\Re(\Psi_{\mu\nu})\frac{\partial}{\partial K_{\mu}}\frac{\partial}{\partial K_{\nu}}H=0.$ (137)
@@
$\frac{\mathrm{dln}(|A^{(0)}|)}{\mathrm{d}\tau}-\frac{1}{4}\left(\frac{\mathrm{dln}\left(\mathrm{Det(}\Im(\Psi)_{\mu\nu})\right)}{\mathrm{d}\tau}-2\hat{g}_{\mu}\hat{g}_{\nu}(\frac{\partial}{\partial K_{\mu}}\partial_{\nu}H+\Re(\Psi_{\mu k})\partial K_{k}\frac{\partial}{\partial K_{\nu}}H)\right)=0.$ (141)
@@
$\frac{\mathrm{dln}(|A^{(0)}|)}{\mathrm{d}\tau}=\frac{1}{4}\frac{\mathrm{dln}\left(\mathrm{Det(}\Im(\Psi)_{\mu\nu})\right)}{\mathrm{d}\tau}-\frac{1}{2g}\frac{\mathrm{d}g}{\mathrm{d}\tau},$ (142)
@@
$|A^{(0)}|=C\frac{\mathrm{Det(}\Im(\Psi)_{\mu\nu})^{\frac{1}{4}}}{g^{\frac{1}{2}}}.$ (143)
@@
$\begin{split}&\frac{\mathrm{d}\phi}{\mathrm{d}\tau}+\frac{1}{\mathrm{i}}\frac{\mathrm{d}K_{\mu}}{\mathrm{d}\tau}\hat{e}^{*}_{m}\frac{\partial}{\partial K_{\mu}}\hat{e}_{m}+\frac{1}{2}\Im(\Psi_{\mu\nu})\frac{\partial}{\partial K_{\mu}}\frac{\partial}{\partial K_{\nu}}H\\ &+\frac{1}{2\mathrm{i}}\left(\hat{e}^{*}_{i}\frac{\partial}{\partial K_{\mu}}D_{im}\partial_{\mu}\hat{e}_{m}-\partial_{\mu}\hat{e}^{*}_{i}\frac{\partial}{\partial K_{\mu}}D_{im}\hat{e}_{m}\right)=0.\end{split}$ (144)
From here we just use the fact that $\frac{\partial}{\partial K_{\mu}}H\hat{e}_{i}=\frac{\partial}{\partial K_{\mu}}D_{im}\hat{e}_{m}+D_{im}\frac{\partial}{\partial K_{\mu}}\hat{e}_{m}$ along with its conjugate equation to arrive at this final simple expression for the phase evolution:
$\frac{\mathrm{d}\phi}{\mathrm{d}\tau}=\mathrm{i}\hat{e}^{*}_{i}\frac{\mathrm{d}\hat{e}_{i}}{\mathrm{d}\tau}+\frac{\mathrm{i}}{2}\left(\partial_{\mu}\hat{e}^{*}_{i}D_{im}\frac{\partial}{\partial K_{\mu}}\hat{e}_{m}-\frac{\partial}{\partial K_{\mu}}\hat{e}^{*}_{i}D_{im}\partial_{\mu}\hat{e}_{m}\right)-\frac{1}{2}\Im(\Psi_{\mu\nu})\frac{\partial}{\partial K_{\mu}}\frac{\partial}{\partial K_{\nu}}H.$ (145)
@@
$\frac{\mathrm{d}\phi_{P}}{\mathrm{d}\tau}=\mathrm{i}\hat{e}^{*}_{i}\frac{\mathrm{d}\hat{e}_{i}}{\mathrm{d}\tau}+\frac{\mathrm{i}}{2}(\partial_{\mu}\hat{e}^{*}_{i}D_{im}\frac{\partial}{\partial K_{\mu}}\hat{e}_{m}-\frac{\partial}{\partial K_{\mu}}\hat{e}^{*}_{i}D_{im}\partial_{\mu}\hat{e}_{m}).$ (146)
@@
$\frac{\mathrm{d}\phi_{G}}{\mathrm{d}\tau}=-\frac{1}{2}\Im(\Psi_{\mu\nu})\frac{\partial}{\partial K_{\mu}}\frac{\partial}{\partial K_{\nu}}H.$ (147)
We have thus successfully defined $A^{(0)}(\tau)$ and $\Psi_{\mu\nu}(\tau)$ such that our terms that do not involve derivatives of our profile function are set to 0.